A wholessale disstribution of a tablet compresion machine finds that the annual demand for the product is normal distribution with mean 140 and standared deviation of 15. ifhe orders only once in a year. What quantity should be ordered to ensure that there is only 5% chance of running short


1
Expert's answer
2021-03-12T07:44:38-0500

Let "X=" the number of units of product: "X\sim N(\mu, \sigma^2)"

Given "\mu=140, \sigma=15."


"P(X>x)=1-P(X\leq x)="

"=1-P(Z\leq \dfrac{x-140}{15})=0.05"

"\dfrac{x-140}{15}=1.645"

"x=165"

165 units should be ordered to ensure that there is only 5% chance of running short.